OpenStudy (anonymous):

Which mixed number falls between 3 1/2 and 5 3/8 ? A. 5 5/12 B. 3 1/4 C. 5 1/3 D. 3 1/6

OpenStudy (campbell_st):

converting the mixed numbers to fractions helps the process you need a number between 3.5 and 5.375 5 5/23 = 5.426666 3 1/4 = 3.25 5 1/3 = 5.3333 3 1/6 = 3.166666 this might make it easier